# Do LLMs Have Distinct and Consistent Personality? TRAIT: Personality Testset designed for LLMs with Psychometrics

### v2

@@ 4. Why not just chill and watch a movie together when you have time?
Figure 15: Word distribution of seeds from ATOMIC10$\times$, used in TRAIT. We extracted the 20 most frequent verbs and the 5 most frequent direct object groups following the method used by Wang et al. (2022).
(a) High Openness Options
(b) Low Openness Options
Figure 16: Word cloud of options in TRAIT-Openness
(a) High Conscientiousness Options
(b) Low Conscientiousness Options
Figure 17: Word cloud of options in TRAIT-Conscientiousness
(a) High Extraversion Options
(b) Low Extraversion Options
Figure 18: Word cloud of options in TRAIT-Extraversion
(a) High Agreeableness Options
(b) Low Agreeableness Options
Figure 19: Word cloud of options in TRAIT-Agreeableness
(a) High Neuroticism Options
(b) Low Neuroticism Options
Figure 20: Word cloud of options in TRAIT-Neuroticism
(a) High Machiavellianism Options
(b) Low Machiavellianism Options
Figure 21: Word cloud of options in TRAIT-Machiavellianism
(a) High Narcissism Options
(b) Low Narcissism Options
Figure 22: Word cloud of options in TRAIT-Narcissism
(a) High Psychopathy Options
(b) Low Psychopathy Options
Figure 23: Word cloud of options in TRAIT-Psychopathy

### v3

@@ 4. Why not just chill and watch a movie together when you have time?
Figure 15: Word distribution of seeds from ATOMIC10$\times$, used in TRAIT. We extracted the 20 most frequent verbs and the 5 most frequent direct object groups following the method used by wang2022self.
@@ A. Yes
(a) High Openness Options
(b) Low Openness Options
(c) High Conscientiousness Options
(d) Low Conscientiousness Options
Figure 20: Word clouds of options in TRAIT-Openness and Conscientiousness
(a) High Extraversion Options
(b) Low Extraversion Options
(c) High Agreeableness Options
(d) Low Agreeableness Options
Figure 21: Word clouds of options in TRAIT-Extraversion and Agreeableness
(a) High Neuroticism Options
(b) Low Neuroticism Options
(c) High Machiavellianism Options
(d) Low Machiavellianism Options
Figure 22: Word clouds of options in TRAIT-Neuroticism and Machiavellianism
(a) High Narcissism Options
(b) Low Narcissism Options
(c) High Psychopathy Options
(d) Low Psychopathy Options
Figure 23: Word clouds of options in TRAIT-Narcissism and Psychopathy